OpenStudy (anonymous):

Divide. Write your answer in simplest form. 3/14 divided by 7/10

OpenStudy (whpalmer4):

Remember when dividing by a fraction, you get the same result by inverting the fraction by which you are dividing, and multiplying instead: \[\frac{\frac{a}{b}}{\frac{c}{d}} = \frac{a}{b}*\frac{d}{c}\]

OpenStudy (anonymous):

i got 15/49.

OpenStudy (anonymous):

thanks